OpenStudy (anonymous):

The ratio of successive amplitudes is A(n+1)=0.8A(n) The amount of energy stored in the oscillation is proportional to the square of the amplitude. Determine for these oscillations, the amount of energy lost as a percentage of the energy stored in the previous oscillation. @experimentX

OpenStudy (experimentx):

what oscillation?? harmonic oscillator?

OpenStudy (anonymous):

Yes, harmonic oscillation.

OpenStudy (anonymous):

This is the graph that's given.

OpenStudy (anonymous):

I did it like this: \[\frac{ E' }{ E } = \frac{ A ^{2} }{ (0.8A)^{2} }\]

OpenStudy (anonymous):

So the energy of the next oscillation would be 0.64 times the energy of previous oscillation. Which means about 36% is lost. Am I right?

OpenStudy (experimentx):

how do you model this recurrence relation? A(n+1)=0.8A(n)

OpenStudy (experimentx):

let your initial condition be A(0)

OpenStudy (anonymous):

Actually, I calculated the ratio of the successive amplitudes. A1/A2=A2/A3=A3/A4 and so on...

OpenStudy (experimentx):

A(n) = 0.8^n A(0)

OpenStudy (anonymous):

How do I do the energy calculation then?

OpenStudy (experimentx):

hmm ... what do you mean by E ??

OpenStudy (experimentx):

KE or ME?

OpenStudy (anonymous):

Energy stored in the oscillation.

OpenStudy (experimentx):

okay okay ...i presume it's ME then ... but BRB in 10 - 15 min

OpenStudy (anonymous):

Okay.

OpenStudy (experimentx):

you know that the equation of harmonic oscillator is given by \[ y = A \sin( \omega t + \phi )\] but A is dependent on ,,, time period.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

How does this relate with energy of oscillation?

OpenStudy (experimentx):

calculate PE from y calculate KE form y'

OpenStudy (experimentx):

but you need to model \( A = A(t) \) such that \( A(n+1) = 0.8 A(n) \), we know that I comes back to same after time T, therefor it would be wise to model the system by using \[ \huge y(t) = 0.8 ^{\frac{\omega t}{2\pi}} A(0) \sin( \omega t) \]

OpenStudy (experimentx):

at some time \( \tau \) find the PE and KE. to find KE, you will need \( y'(x) \) so \[ KE = \frac m 2 y'(x)^2\]

OpenStudy (experimentx):

to find PE you need hooks law ... use \[ PE = \frac 1 2 ky(x)^2 \]

OpenStudy (experimentx):

find the total energy at time \( \tau \) then find total energy at \( \tau + T \) and see what you get and what you an do.

OpenStudy (experimentx):

Damn!! I only slept for 1 hrs in past 50 hrs.

OpenStudy (anonymous):

I'm so sorry. But I think you are just complicating things.

OpenStudy (experimentx):

what made things complicated?

OpenStudy (anonymous):

See, its just a 2-mark question. The graph simply shows an object undergoing SHM that is underdamped. I calculated that the ratio of successive amplitudes stays constant i.e every successive amplitude is 0.8 times that of the previous amplitude. The question states that energy is proportional to the square of amplitude, and asks to calculate how much energy would be lost in each successive amplitude. Isn't it a simple ratio question?

OpenStudy (experimentx):

of course ... if you know the Energy is proportional to square of amplitudes then why not do it yourself lol??

OpenStudy (anonymous):

I asked you if I did it correctly or not. I calculated the square of 0.8 which is 0.64. So I said that energy loss would be 36%. But the answer apparently is 64% loss. Isn't 64% the energy stored in the next oscillation? Which would mean that 36% has been lost?

OpenStudy (experimentx):

i think 64% should be stored as well ... check your question again.

OpenStudy (anonymous):

So you mean that 36% would be lost, right?

OpenStudy (experimentx):

yep ... of course.

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (experimentx):

I am too sleepy right now ... i'll see this in detail tmr.

OpenStudy (anonymous):

Okay. (: